# Biomedical Open Source Software: Crucial Packages and Hidden Heroes

### v2

@@ Abstract
Despite the importance of scientific software for research, it is often not formally recognized and rewarded. This is especially true for foundation libraries, which are used by the software packages visible to the users, being “hidden” themselves. Research stakeholders like funders, infrastructure providers and other organizations need to understand the complex network of computer programs that the modern research relies upon.
In this work, we used the CZ Software Mentions Dataset to map the dependencies of the software used in biomedical papers and find the packages critical to the software ecosystems. We propose centrality metrics for the network of software dependencies, analyze three ecosystems (PyPi, CRAN, Bioconductor) and determine the packages with the highest centrality.
@@ 1 Introduction
Since the second half of the last century, a computer has become as ubiquitous a tool of a scientific lab as balance scales and Bunsen burner in the previous ages. As a consequence, computer software is now crucial to research, bringing new methods and new scale, while offering new potential for reproducibility and extension. This is true not only for the sciences, but for scholarship more broadly (e.g., sociology and the humanities), making the software revolution both wide and deep. Yet we have very limited insight into the software actually used in research. This lack of infrastructural understanding means we are limited in our ability to reward developers and maintainers, encourage collaboration and coordination, and to direct science funding in a well-informed manner.
This problem is exacerbated by the fact that today, software is a complex ecosystem with a myriad of dependencies. When a researcher uses a program, they rely on the work of the people that wrote, tested and maintained it—as well as the people that wrote, tested and maintained the packages and libraries that the program depends on, either at runtime or during the compilation phase. Thus, funders and others who want to support and accelerate science need to find the critically important pieces, both among the user-facing programs and those that are not known to the end users, i.e., the former’s dependencies. This complex nature of the network of dependencies has a number of implications, including for security (Goodin, 2024) and computational reproducibility (Samuel and Mietchen, 2024).
@@
Scientific software is often invisible in publications, because citation practices in science have not changed at the same pace that software has become crucial. For example, software is infrequently and inconsistently formally cited (Howison and Herbsleb, 2011; Howison et al., 2015; Singh Chawla, 2016; Howison and Bullard, 2016; Knowles et al., 2021; Druskat et al., 2024). There have been recent efforts to extract informal citations from the full text of articles (Schindler et al., 2021; Du et al., 2021; Istrate et al., 2022b, a) and to evaluate the “importance” of software packages by looking at papers that cite them (Bogart et al., 2020). Unfortunately, publications sometimes do not mention all the software used in the course of the research.
Besides this, there is another kind of invisibility of scientific software. The programs visible to the user may rely on many other software packages (known in the software world as dependencies). While the end users may mention a package at the top of the dependencies stack, they are likely not even aware of the packages that are further below. These may be packages that the used program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may – and often do – depend on further packages. These latter packages thus become indirect, transitive dependencies of the user-facing program. All the work undertaken to develop, maintain, test, and distribute the underlying software is therefore not visible in the publication record itself, and is not included in derivatives such as citation graphs and knowledge graphs. The situation resembles the famous XKCD cartoon (Munroe, 2020), where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003”. The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs. There have been proposals to assign credit to these packages (Katz, 2014; Katz and Smith, 2014), but so far, the problem has not been solved. To implement the proposed transitive credit mechanisms, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc. One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) (Druskat et al., 2021). Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation graphs (Druskat, 2020). At present, the situation is quite different: sometimes even the maintainers of the down-stack software are not aware of the top stack programs that depend on their work. They need this knowledge when making breaking changes to their packages, which might negatively influence the software that depends on them (Bogart et al., 2021).
@@
By classifying software packages as visible to the end users or primarily important for other packages, we follow the ideas described by Donald E. Stokes in his famous book (Stokes, 1997). There, Stokes distinguishes between applied research with the results visible to the general public, and the pure research, which is less visible, but important for the applied research. Following his ideas, we can put the software packages into a two dimensional plane with the axes corresponding to the frequency of software mentions and a measure of network centrality (Figure 1). The majority of the packages will occupy the lower left corner of the plot, having a fair usage both by the authors of scientific papers and the authors of the software package. These groups may intersect. Popular packages are used by many authors. We are interested in the “Nebraska” packages, not very visible, but of critical importance. Of course, one can think about highly visible and critical “Pasteur” packages, which are used both directly and as a foundation for other libraries.
Figure 1: Classification of software packages inspired by (Stokes, 1997)
@@
We had a fortunate opportunity to explore these ideas due to the generosity of the Chan Zuckerberg Initiative (CZI). On 24–27 October 2023, CZI hosted a hackathon on Mapping the Impact of Research Software in Science (see https://github.com/chanzuckerberg/software-impact-hackathon-2023). One of the projects at this hackathon was Tracing the dependencies of open source software mentioned in the biomedical literature (Brown et al., 2023). In this project, we explored the dependencies of the open source software packages mentioned in the CZ Software Mentions dataset (Istrate et al., 2022b). We decided to limit the study to open source packages: first, because the dependencies of closed source packages are not public, and second, because we believe in the importance of open source for open and reproducible science. The result of this project is this paper. Here, we explore approaches to making the software package infrastructure underlying science more visible. We examine our findings to develop questions to better understand the idea of criticality and opportunities for improving science through adjustments to the research software ecosystems.
@@ 2.1 Network construction
We combined two datasets: The CZ Software Mentions dataset (Istrate et al., 2022b) (which was built using a model trained on the Softcite Dataset gold standard dataset (Du et al., 2021)), and the Ecosyste.ms software dependency dataset (Nesbitt, 2023), which is built by gathering and normalizing the dependency information from three package managers, one for Python (PyPI) and two for R (CRAN and Bioconductor). The authorship data was scraped from Github metadata, for want of consistently available authoritative software authorship metadata.
The CZ Software Mentions dataset (Istrate et al., 2022b, a) identifies which mentions were traced to which ecosystem. In this way, package names from CRAN, Bioconductor and PyPi that have been parsed by the software mentions project are collected. We used the latest release of the package at time of data collection (October 2023) along with the names of their dependencies. Those dependency names have then also been recursively fetched using the most recent release and dependencies until the full list of transitive dependencies was obtained. Limitations of this dependency resolution approach are discussed below.
After this data processing, we produced a two-mode network, with nodes for papers and nodes for packages. Edges from papers to packages were added when a package was mentioned in the full-text of the paper. Edges between packages were added when package manager descriptions indicated a dependency (including both required and suggested dependencies, and including packaging related dependencies such as those needed for development, as well as testing and building documentation). The full code for the processing is available at (Brown et al., 2023). The network is available at (Brown, 2023) in GEXF format (GEXF Working Group, 2009). It has four classes of nodes:
@@
papers from the CZ Software Mentions dataset. They are identified by their Digital Object identifier (DOI). To estimate the impact of the papers, we separately downloaded their citation numbers from OpenAlex (Priem et al., 2022) as of November 2023 (a copy is available in the data subdirectory of (Brown et al., 2023)).
@@
Software from the corresponding ecosystems. We used CZI ID ((Istrate et al., 2022b, a) as the identifier. We did not attempt to identify the same software across the ecosystems (see the discussion below).
The edges are directed and weighted. An edge from software node $A$ to software node $B$ means that software $A$ depends on software $B$ as determined by the corresponding package manager. An edge from a paper node to a software node means that the given software is mentioned in the paper, the weight corresponds to the number of citations the paper received. In what follows, we will consider both the weighted and unweighted version of this network, since the weights have pros and cons, as discussed below.
@@ 2.2 Network analysis
To analyze the network, we relied on directed centrality analysis (Borgatti and Everett, 2021). There are several possible options for centrality measures. Our choice was determined by the following considerations. First, we wanted a centrality metric that can account for papers even though they have no incoming edges in our network (in other words, papers do not receive centrality but should contribute to the centrality of software packages). This criterion excludes eigenvector centrality (Bonacich, 2007). Second, we wanted a centrality metric that can account for the weight of edges from papers to software packages (i.e., some papers contribute more centrality through their citation counts). This criterion excludes the PageRank algorithm (Brin and Page, 1998) which normalizes weights of out-degree. We found that Katz centrality (Katz, 1953) satisfies all our criteria. This approach gives us the attenuation factor $\beta$ as a free parameter to control the importance of papers (which we set equal to 1), such that paper nodes contribute a factor proportional to $\beta$ times their citation count to the software packages they mention. In turn, these software packages contribute a factor proportional to their own centrality to their dependencies. A package can therefore be central by receiving mentions from well-cited papers, by having central dependents, or by a combination of both.
@@
Figure 2: (a) Network visualization of software packages from three ecosystems (from CRAN in green, PyPI in blue, and Bioconductor in pink) connected through their dependencies within their ecosystem and interconnected through papers that mention them. We label the top 3 most central packages in each ecosystem: ggplot2, SAM, and PRISMA for CRAN, velvet, tophat and pymol for PyPI and DeSeq2, edgeR and limma for Bioconductor. The core of the network is dominated by CRAN and PyPI dependencies, despite the fact that three of the five most central packages come from Bioconductor. (b) The top part of the above network, with papers added (in grey) to illustrate how PRISMA can be central due to many mentions in papers.
@@ 3 Results
In Figure 2, we show the overall network of software packages connected through their dependencies within their ecosystem and interconnected through the papers that mention them. Edges from papers to software packages are directed and weighted by the number of citations the paper received. Edges between software dependencies are directed from a dependent to a dependency.
@@
We found a dense core of popular packages that receive many mentions (e.g. ggplot2 in CRAN, tophat in PyPI and limma in Bioconductor), some of which have many dependencies themselves (e.g. ggplot2). We also found less central packages mentioned heavily by specific communities (e.g. PRISMA in CRAN or pymol in PyPI).
@@
Analyzing the entire ecosystems of software dependency, we found that roughly 10% of software packages are part of dependency loops (i.e., cycles in the dependency networks). Interestingly, we found no cycles whatsoever in the connected component of networks that have received software mentions, and their dependencies. We discuss this finding below.
@@
In Figure 3, we show the distribution of packages from the three investigated ecosystems over Katz centrality and mention counts. We used three different ways to calculate centrality: (a) unweighted graph, (b) weighted complete graph, and (c) the largest connected component of the weighted graph. It is remarkable that all three models give rather similar results for many packages. In the next section, we discuss the insights from the cases when the different methods do not agree. As expected, we found a dense cluster of packages with low Katz centrality and low mention counts, i.e., the “Majority” quadrant (see Figure 1).

### v3

@@ Abstract
Despite the importance of scientific software for research, it is often not formally recognized and rewarded. This is especially true for foundational libraries, which are hidden below packages visible to the users (and thus doubly hidden, since even the packages directly used in research are frequently not visible in the paper). Research stakeholders like funders, infrastructure providers, and other organizations need to understand the complex network of computer programs that contemporary research relies upon.
In this work, we use the CZ Software Mentions Dataset to map the upstream dependencies of software used in biomedical papers and find the packages critical to scientific software ecosystems. We propose centrality metrics for the network of software dependencies, analyze three ecosystems (PyPi, CRAN, Bioconductor), and determine the packages with the highest centrality.
Author summary
Scientists today use software as an important tool for their research. The progress of science depends on the people who write software being properly recognized, rewarded, and incentivised. This means that those who fund research or promote scientists need to know the impact of software and identify the most important programs.
The universe of scientific software is complex: some people write programs used by researchers, while some people write programs used by other programs. The latter are often even less visible than the former, but they are important for the progress of science.
In this work, we establish metrics to measure the impact of both types of software. We introduce the network of dependencies (program A used by researchers depends on programs B, C, and D) and use it to calculate the impact of a large number of packages used in the biomedical literature. In doing so, we identify important software packages, regardless of their direct visibility to end users.
We conclude with a discussion of the limitations of our methodology and approaches to improve network validity.
@@ 1 Introduction
Since the second half of the last century, a computer has become as ubiquitous a tool of a scientific lab as balance scales and a Bunsen burner were in previous ages. As a consequence, computer software is now crucial to research, bringing new methods and new scale, while offering new potential for reproducibility and extension. This is true not only for the natural sciences and mathematics but for scholarship more broadly (e.g., sociology and the humanities), making the software revolution both wide and deep. Yet we have very limited insight into the software actually used in research. This lack of infrastructural understanding means we are limited in our ability to reward developers and maintainers, encourage collaboration and coordination, and direct science funding in a well-informed manner.
Fig 1: XKCD on software dependency and ecosystem fragility [1]. Also has roll-over text of Someday ImageMagick will finally break for good and we’ll have a long period of scrambling as we try to reassemble civilization from the rubble. Used with permission: https://creativecommons.org/licenses/by-nc/2.5/
@@
Scientific software is often invisible in publications, because citation practices in science have not changed at the same pace as software has become crucial. For example, software is infrequently and inconsistently formally cited [2, 3, 4, 5, 6, 7]. There have been recent efforts to extract informal citations from the full text of articles [8, 9, 10, 11] and to evaluate the “importance” of software packages by looking at papers that cite them [12]. Unfortunately, publications sometimes do not mention all the software used in the course of research.
Besides this, there is another kind of invisibility of scientific software. The programs visible to the end user may rely on many other software packages (known in the software world as dependencies). While the end users may mention a package at the top of the dependency stack, they are likely not even aware of the full set of packages that are further below. These may be packages that the utilized program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may depend on further packages. These latter packages thus become indirect, transitive dependencies of the user-facing program. This complex nature of the network of dependencies has a number of implications, including for security [13] and computational reproducibility [14]. In particular, all the work undertaken to develop, maintain, test, and distribute the underlying software is not directly visible in the publication record itself, and is not included in derivatives such as citation networks and knowledge graphs.
The situation resembles the famous XKCD cartoon shown in  Figure 1, where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [1]. The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs. The absence of recognition may lead to dire consequences, for example, if, due to the lack of funding, the underlying libraries become vectors for malware attacks.
There have been proposals to assign credit to these packages using the dependency structure to calculate what was been termedtransitive credit [15, 16] but so far the problem has not yet been solved. To implement the proposed measure, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc. One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) [17]. Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation networks [18].
At present, though, the situation is quite different: sometimes even the maintainers of the lower-stack software are not aware of the upper-stack programs that depend on their work. They need this knowledge when making breaking changes to their packages, which might negatively influence the software that depends on them [19].
By classifying software packages as visible to the end users or primarily important for other packages, we follow the ideas described by Donald E. Stokes in his famous book, Pasteur’s Quadrant [20]. There, Stokes distinguishes between applied research with the results visible to the general public, and pure research, which is less visible, but important for the applied research. He discussed a two-dimensional diagram of scientific works, with the importance of the works for the applications on the $x$ axis and the quest for fundamental understanding on the $y$ axis. Thus, the upper left quadrant is occupied by the purely fundamental research personified by Niels Bohr. The lower right quadrant is occupied by purely applied research personified by Thomas Edison. The upper right quadrant is occupied by the works that have both fundamental and practical value. Stokes chose Louis Pasteur as an example of such works. While this picture is rather simplified (Bohr’s works have significant practical value, and Edison’s experiments stimulated fundamental research), it gives important insight into scientific endeavor.
@@
Following these ideas, we can put the software packages into a two-dimensional plane with the axes corresponding to the frequency of software mentions and a measure of network centrality (Figure 2). The majority of the packages will likely occupy the lower left corner of the plot, having a small number of mentions from the authors of scientific papers and limited centrality from usage in other software packages. Popular packages are used by many authors. We are interested in the “Nebraska” packages, not very visible, but of critical importance. Of course, one can think about highly visible and critical “Pasteur” packages, which are used both directly and as a foundation for other libraries.
Fig 2: Classification of software packages inspired by [20]
We had a fortunate opportunity to explore these ideas due to the generosity of the Chan Zuckerberg Initiative (CZI). On 24–27 October 2023, CZI hosted a hackathon on Mapping the Impact of Research Software in Science (see https://github.com/chanzuckerberg/software-impact-hackathon-2023). One of the projects at this hackathon was Tracing the dependencies of open source software mentioned in the biomedical literature [21]. In this project, we explored the dependencies of the open source software packages mentioned in the CZ Software Mentions dataset [10]. The dataset was created using a machine learning system trained on the SoftCite gold standard dataset of manually annotated software mentions [9] and consists of extractions from 2.4 million biomedical papers. The methodology and its limitations are discussed in [11]. In particular, the work on the disambiguation of packages can be improved. Later, we discuss how these limitations influence our results.
@@
We decided to limit our study to open-source packages: first, because the dependencies of closed-source packages are not public, and second, because we believe in the importance of open source for open and reproducible science. Here, we explore approaches to making the software package infrastructure underlying science more visible. We examine our findings to develop questions to better understand the idea of criticality and opportunities for improving science through adjustments to the research software ecosystems.
@@ 2.1 Network construction
We combined three datasets: The CZ Software Mentions dataset [10], which was constructed using a machine learning model trained on the SoftCite Gold Standard annotations [9], the Ecosyste.ms software dependency dataset [22], which is built by gathering and normalizing the dependency information from multiple software ecosystems, including one for the Python programming language (PyPI) and two for the R programming language (CRAN and Bioconductor), and finally we used OpenAlex [23] for information on how frequently papers had been cited. The software authorship data were scraped from GitHub metadata.
The CZ Software Mentions dataset [10, 11] includes mentions from 2.4 million biomedical papers and identifies which mentions were traced to which ecosystem. In this way, package names from CRAN, Bioconductor, and PyPI are parsed by the software mentions extraction model, collected, and finally disambiguated. We collected information for each package, including its dependencies, using the latest release of the package at the time of data collection (October 2023). Dependencies were then recursively retrieved using the most recent release and following dependencies until the full list of transitive dependencies was obtained. Limitations of our dependency resolution approach are discussed later.
After this data processing, we produced a two-mode network, with nodes for papers and nodes for software packages. Edges from papers to software packages were added when a package was mentioned in the full text of the article. Edges between packages were added when metadata descriptions indicated a required dependency. The full code for the processing is available at [21]. The network is available at [24] in GEXF format [25] and has four classes of nodes:
@@
papers from the CZ Software Mentions dataset. They are identified by their Digital Object Identifier (DOI). To estimate the impact of the papers, we separately downloaded their citation numbers from OpenAlex [23] as of November 2023 (a copy is available in the data subdirectory of [21]).
@@
Software from the corresponding ecosystems. We used CZI ID ([10, 11] as the identifier. We did not attempt to identify the same software across the ecosystems (see the discussion below).
The edges are directed and weighted. An edge from one software node $A$ to another software node $B$ means that software $A$ depends on software $B$ as determined by the corresponding metadata. An edge from a paper node to a software node means that the given software is mentioned in the paper, and the weight corresponds to the number of citations the paper received. In what follows, we will consider both the weighted and unweighted versions of this network, as utilization of the weights has both positive and negative aspects to consider during analysis and interpretation.
@@ 2.2 Network analysis
To analyze the network, we relied on directed centrality analysis [26]. There are several possible options for centrality measures. Our choice was determined by the following considerations. First, we wanted a centrality metric that can account for papers even if they have no incoming edges in our network (in other words, papers do not receive centrality but should contribute to the centrality of software packages). This criterion excludes eigenvector centrality [27]. Second, we wanted a centrality metric that can account for the weight of edges from papers to software packages (i.e., some papers contribute more centrality because they are more highly cited.). This criterion excludes the PageRank algorithm [28], which normalizes weights of out-degree. We found that Katz centrality [29] satisfies all our criteria. Katz centrality can be interpreted as the importance of the node for the diffusion processes on the network. This corresponds to the intuition that innovation is a diffusion-like process [30]. Katz centrality gives us the attenuation factor $\beta$ as a free parameter to control the importance of papers (which we set equal to 1), such that paper nodes contribute a factor proportional to $\beta$ times their citation count to the software packages they mention (we follow the formulation of Katz centrality in [31]). In turn, software packages contribute a factor proportional of their own centrality to their dependencies. A package can therefore be central by receiving mentions from well-cited papers, by having central dependents, or by a combination of both.
@@
Fig 3: (a) Network visualization of software packages from three ecosystems (from CRAN in green, PyPI in blue, and Bioconductor in pink) connected through their dependencies within their ecosystem and interconnected through papers that mention them. We label the top 3 most central packages in each ecosystem: ggplot2, SAM, and PRISMA for CRAN, velvet, tophat and pymol for PyPI and DeSeq2, edgeR and limma for Bioconductor. The core of the network is dominated by CRAN and PyPI dependencies, despite the fact that three of the five most central packages come from Bioconductor. (b) The top part of the above network, with papers added (in grey) to illustrate how PRISMA can be central due to many mentions in papers.
@@ 3 Results
In Figure 3, we show the overall network of software packages connected by their dependencies within each ecosystem and interconnected through the papers that mention them. Edges from papers to software packages are directed and weighted by the number of citations the paper received. Edges between software dependencies are directed from a dependent to a dependency.
@@
We found a dense core of popular packages that receive many mentions (e.g. ggplot2 in CRAN, tophat in PyPI and limma in Bioconductor), some of which have many dependencies themselves (e.g. ggplot2). We also found packages specific to certain communities (e.g. PRISMA in CRAN or pymol in PyPI).
@@
Analyzing the entire ecosystem of software dependencies, we found that roughly 10% of software packages are part of dependency loops (i.e., cycles in the dependency networks). Interestingly, we found no cycles in the connected components of networks that have received software mentions, and their dependencies. We discuss this finding below. Note that contemporary package managers usually resolve the dependency loops by installing all packages in the loop simultaneously, so the loops, while indicating rather sloppy software practices, are not fatal.
@@
In Figure 4, we show the distribution of packages from the three investigated ecosystems over Katz centrality and mention counts. We used three different ways to calculate centrality: (a) unweighted graph, (b) weighted complete graph, and (c) the largest connected components of the weighted graph (calculated separately for each ecosystem). It is remarkable that all three models give rather similar results for many packages. In the next section, we discuss the insights from the cases when the different methods do not agree. As expected, we found a dense cluster of packages with low Katz centrality and low mention counts, i.e., the “Majority” quadrant (see Figure 2).
Table 1: Summary statistics for the number of mentions in papers. “Number” is the number of packages, “Dependency only” is the fraction of packages not mentioned in any papers and pulled by the dependencies only, “Median” is the median number of mentions, “IQR” is the interquartile range for mentions count, “Gini” is Gini coefficient for the mentions count.
Ecosystem Number Dependency only Mentions count
Median IQR Max Gini
bioconductor $1018$ 0.047 151 277 013 752 50.04715127701375250.047\,151\,277\,013\,752\,50.047 151 277 013 752 5 $24$ $68$ 21 2572125721\,25721 257 0.843 712 143 943 3090.8437121439433090.843\,712\,143\,943\,3090.843 712 143 943 309
cran $3594$ 0.113 244 296 048 970.113244296048970.113\,244\,296\,048\,970.113 244 296 048 97 $9$ $34$ 21 9602196021\,96021 960 0.889 196 914 901 7220.8891969149017220.889\,196\,914\,901\,7220.889 196 914 901 722
pypi $5596$ 0.143 495 353 824 160.143495353824160.143\,495\,353\,824\,160.143 495 353 824 16 $4$ $17$ 20 3222032220\,32220 322 0.888 156 736 619 3960.8881567366193960.888\,156\,736\,619\,3960.888 156 736 619 396
Table 2: Summary statistics for Katz centrality. “Median” is the median centrality, “IQR” is the interquartile range, “Max” is the maximal centrality, “Gini” is Gini coefficient.
Ecosystem Median IQR Max Gini
Unweighted graph
bioconductor 0.001 263 667 285 359 560.001263667285359560.001\,263\,667\,285\,359\,560.001 263 667 285 359 56 0.000 651 274 677 839 160.000651274677839160.000\,651\,274\,677\,839\,160.000 651 274 677 839 16 0.217 934 004 136 6260.2179340041366260.217\,934\,004\,136\,6260.217 934 004 136 626 0.549 779 090 129 2210.5497790901292210.549\,779\,090\,129\,2210.549 779 090 129 221
cran 0.001 166 462 109 562 670.001166462109562670.001\,166\,462\,109\,562\,670.001 166 462 109 562 67 0.000 486 025 878 984 450.000486025878984450.000\,486\,025\,878\,984\,450.000 486 025 878 984 45 0.277 617 982 075 9160.2776179820759160.277\,617\,982\,075\,9160.277 617 982 075 916 0.496 401 393 691 9490.4964013936919490.496\,401\,393\,691\,9490.496 401 393 691 949
pypi 0.001 080 921 554 861 410.001080921554861410.001\,080\,921\,554\,861\,410.001 080 921 554 861 41 0.000 194 410 351 593 780.000194410351593780.000\,194\,410\,351\,593\,780.000 194 410 351 593 78 0.191 008 170 440 8880.1910081704408880.191\,008\,170\,440\,8880.191 008 170 440 888 0.351 314 534 735 6710.3513145347356710.351\,314\,534\,735\,6710.351 314 534 735 671
Weighted, complete graph
bioconductor 0.001 431 553 182 867 530.001431553182867530.001\,431\,553\,182\,867\,530.001 431 553 182 867 53 0.000 322 302 811 767 7570.0003223028117677570.000\,322\,302\,811\,767\,7570.000 322 302 811 767 757 0.105 238 137 865 5360.1052381378655360.105\,238\,137\,865\,5360.105 238 137 865 536 0.358 203 102 844 1960.3582031028441960.358\,203\,102\,844\,1960.358 203 102 844 196
cran 0.001 388 172 783 386 690.001388172783386690.001\,388\,172\,783\,386\,690.001 388 172 783 386 69 0.000 147 428 626 544 9850.0001474286265449850.000\,147\,428\,626\,544\,9850.000 147 428 626 544 985 0.128 618 817 548 2220.1286188175482220.128\,618\,817\,548\,2220.128 618 817 548 222 0.253 046 882 585 3520.2530468825853520.253\,046\,882\,585\,3520.253 046 882 585 352
pypi 0.001 382 750 233 451 590.001382750233451590.001\,382\,750\,233\,451\,590.001 382 750 233 451 59 0.000 104 384 086 250 760.000104384086250760.000\,104\,384\,086\,250\,760.000 104 384 086 250 76 0.073 772 436 229 610.073772436229610.073\,772\,436\,229\,610.073 772 436 229 61 0.180 193 304 141 0160.1801933041410160.180\,193\,304\,141\,0160.180 193 304 141 016
Weighted, LCC only
bioconductor 0.002 883 015 709 935 830.002883015709935830.002\,883\,015\,709\,935\,830.002 883 015 709 935 83 0.000 759 200 923 709 920.000759200923709920.000\,759\,200\,923\,709\,920.000 759 200 923 709 92 0.210 742 476 047 3810.2107424760473810.210\,742\,476\,047\,3810.210 742 476 047 381 0.371 663 779 294 7510.3716637792947510.371\,663\,779\,294\,7510.371 663 779 294 751
cran 0.002 793 430 476 011 270.002793430476011270.002\,793\,430\,476\,011\,270.002 793 430 476 011 27 0.000 369 917 791 154 9370.0003699177911549370.000\,369\,917\,791\,154\,9370.000 369 917 791 154 937 0.257 562 976 941 2260.2575629769412260.257\,562\,976\,941\,2260.257 562 976 941 226 0.280 390 754 327 0070.2803907543270070.280\,390\,754\,327\,0070.280 390 754 327 007
pypi 0.002 785 286 363 836 310.002785286363836310.002\,785\,286\,363\,836\,310.002 785 286 363 836 31 0.000 312 869 642 721 380.000312869642721380.000\,312\,869\,642\,721\,380.000 312 869 642 721 38 0.147 731 480 149 7160.1477314801497160.147\,731\,480\,149\,7160.147 731 480 149 716 0.220 244 615 942 4560.2202446159424560.220\,244\,615\,942\,4560.220 244 615 942 456
Figure 4 suggests large differences between the packages in the majority quadrant and the packages in the rest of the system. This is further illustrated by Tables1 and 2. The sum of the median and half of the interquartile range is several orders of magnitude smaller than the maximal value for both the number of mentions and centrality.
A convenient measure of this inequality is the Gini coefficient [32]. Gini coefficients close to one characterize a very unequal distribution, while Gini coefficients close to zero correspond to a uniform distribution. As seen from Table 1, the distribution of mention counts is wildly unequal with high Gini coefficients. The distribution of centrality (Table 2) is smoother, especially for the weighted graph, but is still rather unequal.
Perhaps an even more robust measure of “Nebraska” properties of the ecosystem is the number of packages that are not mentioned in any paper, but are in the dependencies of the packages that are mentioned (we are grateful to the anonymous reviewer of PLoS CompBio for suggesting this metric). It is shown in Table 1.
In Tables 3, 4, and 5 we show the packages in “Pasteur”, Popular, and “Nebraska” packages using the following criteria:
  1. 1.
    “Pasteur”: the highest sum of mentions and centrality percentile.
  2. 2.
    Popular: the highest mentions percentile while centrality percentile is less than $0.5$.
  3. 3.
    “Nebraska”: the highest centrality percentile while mentions percentile is less than $0.5$.